OpenStudy (anonymous):

evaluate lim t approaches -3 (t^2-9)/(2t^2+7x+3)

OpenStudy (anonymous):

You need to use l'Hopitals rule for this one.

OpenStudy (anonymous):

12/5

OpenStudy (anonymous):

yes i know i factor out the numerator now i need help on the deonomotinor

OpenStudy (anonymous):

Oh. I suppose factoring works well here.. \[2t^2 + 7t + 3 = (2t + 1)(t + 3)\] So t+3 will cancel with the t+3 on the top.

OpenStudy (anonymous):

And you get (-3-3)/(2(-3)+1) = -6/-5 = 6/5

OpenStudy (anonymous):

thanks again